# Representing Molecules as Random Walks Over Interpretable Grammars

### v1

@@ Abstract
Recent research in molecular discovery has primarily been devoted to small, drug-like molecules, leaving many similarly important applications in material design without adequate technology. These applications often rely on more complex molecular structures with fewer examples that are carefully designed using known substructures. We propose a data-efficient and interpretable model for representing and reasoning over such molecules in terms of graph grammars that explicitly describe the hierarchical design space featuring motifs to be the design basis. We present a novel representation in the form of random walks over the design space, which facilitates both molecule generation and property prediction. We demonstrate clear advantages over existing methods in terms of performance, efficiency, and synthesizability of predicted molecules, and we provide detailed insights into the method’s chemical interpretability.
@@ 1 Introduction
In terms of quality, we demonstrate our grammar representation suits applications in materials science. We select datasets that reflect real-world settings of experimentally curated designs of molecules with complex, modular sub-structures characterized by functional groups known or hypothesized to yield high target properties.
@@ 2 Related Works
Recent work has shown that such grammars represent a data-efficient way for representing molecules and yield SOTA results (Guo et al., 2023a, b). In a nutshell, this is achieved in that (Geo-)DEG model by explicitly representing the training data’s design space in terms of learnt motifs, in the form of a graph grammar. Grammars naturally allow for generating novel molecules in the given design space. Yet, obtaining production rules involves either manual definition (Krenn et al., 2020; Guo et al., 2022; Nigam et al., 2021) or a significant complexity to automatically learn (Guo et al., 2023a; Kajino, 2019), where the training times for downstream tasks are considerable (see Figure 5). Further, the learnt substructures sometimes lack a chemical interpretation, and grammar derivations often produce in chemically invalid structures (Guo et al., 2023a), so the natural potential of symbolic methods for interpretability and validity is lost, although such elements are critical for expert validation and for gaining scientific insights. We propose a novel way for representing and learning such context-sensitive grammars, over a design space informed by chemical motifs. This approach results in order-of-magnitude differences in runtime and enhances chemical interpretability.
@@
There are various other non-motif based approaches that we compare to in our evaluation, namely, (pre-trained) GNNs (Hu et al., 2020; Xia et al., 2023b), motif-based pre-training approaches designed for semi- or unsupervised learning (Xia et al., 2023b), and molecular few-shot learning including the SOTA, which relies on modeling the domain expert’s reasoning in terms of related molecule contexts using associative memories (Schimunek et al., 2022). Central to our method is the connection between random walks and graph diffusion, established methods that have been shown to be particularly effective to model graph structures through physics-inspired processes (Thanou et al., 2017). Other related works and more detailed discussion can be found in Appendix C.
@@ 3 An Interpretable, Grammar-based Molecule Representation and Efficient Learning
Figure 1: Illustration of our random walk representation: (a) (top) molecule $M$, number 33 (middle) $H_{M}$ as a connected subgraph of $G$ (bottom) $\hat{H}_{M}$ as a random walk over $H_{M}$; (b) the motif graph $G$, each node is a motif $v$ that contains both the molecular fragment $v_{B}$ (black molecule sections) and the contexts for attachment ($v_{R}$, red molecule sections), each gray line indicates a possible attachment between nodes. Directed edges of $\hat{H}_{M}$ use the same color as the dashed border of the corresponding figure of $M$; (c) (top) demonstration of motif matching criteria eq 1-4 ($183\leftrightarrow 5$) (bottom) two more examples of $H_{M}$.
Our method employs a graph grammar, which is composed of a set of predefined molecular motifs and a set of transition rules. The motifs are devised either through automatic generation or manual curation and are interconnected following the transition rules to assemble into a complete molecular structure. Following (Guo et al., 2023a, b), a grammar $\mathcal{G}\mkern 3.0mu{=}\mkern 3.0mu(\mathcal{N},\Sigma,\mathcal{P},\mathcal{X})$ contains a set $\mathcal{N}$ of non-terminal nodes, a set $\Sigma$ of terminal nodes representing chemical atoms, and a starting node $\mathcal{X}$. The generation of molecular graphs is described using a set of production rules, $\mathcal{P}=\{p_{i}|i=1,\ldots,k\}$. Each rule, $p_{i}$, is defined by a transformation from a left-hand side ($LHS$) to a right-hand side ($RHS$), with both sides being graphs. The process starts from an initial single-node graph $\mathcal{X}$, and a molecule is constructed by iteratively applying a rule from $\mathcal{P}$, where the $LHS$ of the selected rule matches a subgraph within the current graph. This selected subgraph is then replaced by the corresponding $RHS$ of the rule.
We introduce random walk grammar, characterized by a specific condition where the $LHS$ of each rule is strictly contained within its $RHS$. Such a design ensures that the generation of a molecule is a progressive process, where in each step, a new subgraph is attached to the existing graph. We implement the grammar using a compact motif graph $G$ (Fig. 1 (b)); the nodes are the motifs and each edge describes the application of a transition rule. The difference between $RHS$ and $LHS$ of each rule is exactly a motif. We highlight two novelties of this work:
@@
We demonstrate the utility of our grammar-based molecule representation for both molecular generation and property prediction tasks.
@@
Our method builds upon a given molecule fragmentation. More specifically, given a dataset $D=\{M^{(i)}:=(V_{M^{(i)}},E_{M^{(i)}})\}_{i=1}^{|D|}$, a fragmentation of $M^{(i)}$ is a collection of disjoint molecular graphs $\{F^{(i)}_{j}\}:=\{(V^{(i)}_{j},E^{(i)}_{j})\}$ such that $\sqcup_{j}V^{(i)}=V_{M^{(i)}}$. Letting $g(v)$ denote the node-induced subgraph of $g$ by $v$, $F^{(i)}_{j}$ is the subgraph of $M^{(i)}$ induced by $V^{(i)}_{j}$. When $F^{(i)}_{j}$ is a chemical motif, it is essential to know the possible contexts within which $F^{(i)}_{j}$ occur, because the behavior of one substructure is often influenced by neighboring structures [^2] For materials applications that rely in particular on electrophilicity, polarity, and extended aromaticity, longer-range combinations and patterns of motifs are often more influential than any individual one.. Specifically, given neighboring fragments $j_{1},j_{2}$, i.e. $\exists e\in E_{M^{(i)}}\text{ s.t. }e\notin E^{(i)}_{j_{1}},e\notin E^{(i)}_{j_{2}}$ and $e\in M^{(i)}(V^{(i)}_{j_{1}}\cap V^{(i)}_{j_{2}})$, then we can use automatic rules $R_{D}$ to infer the “context” of $j_{1}$: $c^{(j_{1})}_{j_{2}}:=R_{D}(V^{(i)}_{j_{1}},V^{(i)}_{j_{2}})\text{ s.t. }c^{(j_{1})}_{j_{2}}\subseteq V^{(i)}_{j_{2}}$ and $M^{(i)}(V_{j_{1}}^{(i)}\sqcup c^{(j_{1})}_{j_{2}})$ is connected. The descriptions and examples for dataset-specific rules are given in Appendix A.
@@
3.1 The Molecular Design Space as Context-Sensitive Grammar Over Motif Graph
Given a set of motifs, $V$, we describe our hierarchical abstraction over $V$. $G=(V,E)$ is a directed multigraph. Each $v\in V$ contains both the motif graph $g_{v}$ and $\{v_{r_{l}}\}$, denoting the possible “contexts” for attaching $g_{v}$ to another motif; that is, $\forall l,v_{r_{l}}\subseteq N(g_{v})$, with $N(g)$ denoting the set of atom nodes of graph $g$. $v_{R}:=\cup_{l}v_{r_{l}}$, and $\emptyset\neq v_{B}:=N(g_{v})\setminus v_{R}.$ Denoting $\sim$ to be the isomorphism relation, we construct $E$ by matching every pair of motifs $u,v$ and their contexts $(l_{1},l_{2})$ by finding corresponding subgraphs in $u_{B}$ and $v_{B}$ to match $u_{r_{l_{1}}}$ and $v_{r_{l_{2}}}$, as shown in Fig. 1. Specifically, $(u,v,e_{l_{1},l_{2}})\in E\iff\exists b_{2}\subseteq u_{B},b_{1}\subseteq v_{B}$ such that:
@@ 3.2 Molecules as Random Walks in the Design Space
Intuitively, our representation of a molecule $M$ captures a derivation in the above-defined context-sensitive grammar. While prior work has modeled such derivations in large and complex tree structures (e.g., with auxiliary nodes for partial derivations) (Guo et al., 2023a, b), we model it compactly in terms of a random walk over the bidirectionally connected subgraph $H_{M}=(V_{M},E_{M})$ of $G$ given by the fragmentation of $M$ [^3] Refer to Appendix B for how and why we augment $G$ with duplicates of the same motif.; see Figure 1(a). Observe that $G$ is a strong prior for constraining the design space and sufficient for describing the molecular structure of $M$, but $H_{M}$ misses the global distribution of which it is a sample of.
Our learnable component models this distribution and, at the same time, captures the features that characterize a specific molecule in terms of a random walk. More specifically, our final molecule representation is a directed-acyclic multi-graph $\hat{H}_{M}=(V_{M},\hat{E}_{M},w_{M})$ that linearizes $H_{M}$ into a random walk such that (1) $\hat{E}_{M}\subseteq E_{M}$, (2) $\hat{H}_{M}$ remains connected, and (3) there is an Euler path [^4] In the case of monomers, the Euler path needs to be closed as monomers have the property of self-loops.. (i.e., each edge is used exactly once) $v_{0},v_{1},\ldots,v_{\ell}$ over $(V_{M},\hat{E}_{M})$ with $\hat{E}_{M}:=\cup_{i}\{(v_{i},v_{i+1})\}$; this path can be generated via a pre-order traversal that adds a reversed duplicate of the sub-trajectory when the stack contracts. The last component, $w_{M}$ is the sequence $w_{M}:=p_{0},p_{1},\dots,p_{\ell-1}$ of probabilities given by the random walk; that is, $p_{i}$ represents the probability with which the edge between $v_{i}$ and $v_{i+1}$ was traversed in the context of all nodes visited thus far, as shown in Fig. 2. $w_{M}$ is parameterized by our learnable grammar, and $\hat{H}_{M}$ is explainable as a random walk of context-sensitive grammar rule applications.
Figure 2: Illustration of our generation procedure: (t=1) our learnable grammar parameterized by $\Phi$ samples a state transition $56\rightarrow 9$; (t=2) under the context of having visited $56$, our grammar samples a state transition $\rightarrow 71$; (t=10) (bottom) our grammar samples a final transition $5$, which determines the molecular structure (top); our program’s notation is 56971[705]70:15:156$\rightarrow$ 9$\rightarrow$ 71[$\rightarrow$ 70$\rightarrow$ 5$\rightarrow$ 70:1% $\rightarrow$ 5:156 → 9 → 71 [ → 70 → 5 ] → 70 : 1 → 5 : 1
@@ 3.3.1 Parameter Estimation
where $D\in\mathbb{R}^{|V|\times|V|}$ is the in-degree matrix of $G$, $h(\cdot;\phi)$ is a context-sensitive adjustment layer and the learnable parameters are $\Phi=(W,\phi)$. The full training algorithm can be found in Appendix D.
@@ 3.3.2 Training for a Downstream Task
Property Prediction Our grammar-based molecular graph representation $\hat{H}_{i}$ allows for applying an off-the-shelf graph neural network $\mathcal{F}_{\Theta}$ to solve a given prediction task; in our evaluation, we used GIN (Xu et al., 2019). Given a property value $y^{(i)}\in\mathbb{R}$ with each molecule $M^{(i)}$, we apply a linear head $f_{\theta}$ and application-specific loss function $\mathcal{L}$ (e.g., MSE for regression or cross-entropy for classification).

### v2

@@ Abstract
Recent research in molecular discovery has primarily been devoted to small, drug-like molecules, leaving many similarly important applications in material design without adequate technology. These applications often rely on more complex molecular structures with fewer examples that are carefully designed using known substructures. We propose a data-efficient and interpretable model for representing and reasoning over such molecules in terms of graph grammars that explicitly describe the hierarchical design space featuring motifs to be the design basis. We present a novel representation in the form of random walks over the design space, which facilitates both molecule generation and property prediction. We demonstrate clear advantages over existing methods in terms of performance, efficiency, and synthesizability of predicted molecules, and we provide detailed insights into the method’s chemical interpretability. Code is available at https://github.com/shiningsunnyday/polymer_walk.
@@ 1 Introduction
In terms of quality, we demonstrate our grammar representation suits applications characterized by designer molecules. We select datasets that reflect real-world settings of experimentally curated designs of molecules with complex, modular sub-structures characterized by functional groups known or hypothesized to yield high target properties.
@@ 2 Related Works
Recent work has shown that such grammars represent a data-efficient way for representing molecules and yield SOTA results (Guo et al., 2023a, b). In a nutshell, this is achieved by explicitly representing the training data’s design space in terms of learnt motifs, in the form of a graph grammar. Grammars naturally allow for generating novel molecules in the given design space. Yet, obtaining production rules involves either manual definition (Krenn et al., 2020; Guo et al., 2022; Nigam et al., 2021) or a significant complexity to automatically learn (Guo et al., 2023a; Kajino, 2019), where the training times for downstream tasks are considerable (see Figure 5). Further, the learnt substructures sometimes lack a chemical interpretation, and grammar derivations often produce chemically invalid structures (Guo et al., 2023a), so the natural potential of symbolic methods for interpretability and validity is lost, although such elements are critical for expert validation and for gaining scientific insights. We propose a novel way for representing and learning such context-sensitive grammars, over a design space informed by chemical motifs. This approach results in order-of-magnitude differences in runtime and enhances chemical interpretability.
@@
There are various other non-motif based approaches that we compare to in our evaluation, including (pre-trained) GNNs (Hu et al., 2020; Xia et al., 2023b), motif-based pre-training approaches designed for semi- or unsupervised learning (Xia et al., 2023b), and molecular few-shot learning including the SOTA, which relies on modeling the domain expert’s reasoning in terms of related molecule contexts using associative memories (Schimunek et al., 2023). Central to our method is the connection between random walks and graph diffusion, established methods that have been particularly effective to model graph structures through physics-inspired processes (Thanou et al., 2017). Other related works and more detailed discussions can be found in Appendix C.
@@ 3 An Interpretable, Grammar-based Molecule Representation and Efficient Learning
Figure 1: Illustration of our random walk representation: (a) (top) molecule $M$, number 33 (middle) $H_{M}$ as a connected subgraph of $G$ (bottom) $\hat{H}_{M}$ as a random walk over $H_{M}$; (b) the motif graph $G$, each node is a motif $v$ that contains both the molecular fragment $v_{B}$ (black molecule sections) and the contexts for attachment ($v_{R}$, red molecule sections), each gray line indicates a possible attachment between nodes. Directed edges of $\hat{H}_{M}$ use the same color as the dashed border of the corresponding figure of $M$; (c) (top) demonstration of motif matching criteria eq 1-4 ($183\leftrightarrow 5$), another example is in Fig. 11 (bottom) two more examples of $H_{M}$.
Our method employs a graph grammar, which is composed of a set of predefined molecular motifs and a set of transition rules. The motifs are devised either through automatic generation or manual curation and are interconnected following the transition rules to assemble into a complete molecular structure. Following (Guo et al., 2023a, b), a grammar $\mathcal{G}\mkern 3.0mu{=}\mkern 3.0mu(\mathcal{N},\Sigma,\mathcal{P},\mathcal{X})$ contains a set $\mathcal{N}$ of non-terminal nodes, a set $\Sigma$ of terminal nodes representing chemical atoms, and a starting node $\mathcal{X}$. The generation of molecular graphs is described using a set of production rules, $\mathcal{P}=\{p_{i}|i=1,\ldots,k\}$. Each rule, $p_{i}$, is defined by a transformation from a left-hand side ($LHS$) to a right-hand side ($RHS$), with both sides being graphs. The process starts from an initial empty graph $\mathcal{X}$, and a molecule is constructed by iteratively applying a rule from $\mathcal{P}$, where the $LHS$ of the selected rule matches a subgraph within the current graph. This selected subgraph is then replaced by the corresponding $RHS$ of the rule.
Random Walk Grammar.
We introduce random walk grammar, characterized by a specific condition where the $LHS$ of each rule differs from its $RHS$ by exactly a motif. Such a design ensures that the generation of a molecule is a progressive process, where in each step, a new subgraph is attached to the existing graph. We implement the grammar using a compact motif graph $G$ (Fig. 1 (b)); the nodes are the motifs and each edge describes the application of a transition rule.
We highlight two novelties of this work:
@@
We demonstrate the utility of our grammar-based representation for both molecular generation and property prediction tasks. The main steps of our workflow are as follows.
@@
Our method builds upon a given molecule fragmentation. More specifically, given a dataset $D=\{M^{(i)}:=(V_{M^{(i)}},E_{M^{(i)}})\}_{i=1}^{|D|}$, a fragmentation of $M^{(i)}$ is a collection of disjoint molecular graphs $\{F^{(i)}_{j}\}:=\{(V^{(i)}_{j},E^{(i)}_{j})\}$ such that $\sqcup_{j}V^{(i)}=V_{M^{(i)}}$. Letting $g(v)$ denote the node-induced subgraph of $g$ by $v$, $F^{(i)}_{j}$ is the subgraph of $M^{(i)}$ induced by $V^{(i)}_{j}$. When $F^{(i)}_{j}$ is a chemical motif, it is essential to know the possible contexts within which $F^{(i)}_{j}$ occur, because the behavior of one substructure is often influenced by neighboring structures [^2] For materials applications that rely in particular on electrophilicity, polarity, and extended aromaticity, longer-range combinations and patterns of motifs are often more influential than any individual one.. Specifically, given neighboring fragments $j_{1},j_{2}$, i.e. $\exists e\in E_{M^{(i)}}\text{ s.t. }e\notin E^{(i)}_{j_{1}},e\notin E^{(i)}_{j_{2}}$ and $e\in M^{(i)}(V^{(i)}_{j_{1}}\cap V^{(i)}_{j_{2}})$, then we can use automatic rules $R_{D}$ to infer the “context” of $j_{1}$: $c^{(j_{1})}_{j_{2}}:=R_{D}(V^{(i)}_{j_{1}},V^{(i)}_{j_{2}})\text{ s.t. }c^{(j_{1})}_{j_{2}}\subseteq V^{(i)}_{j_{2}}$ and $M^{(i)}(V_{j_{1}}^{(i)}\sqcup c^{(j_{1})}_{j_{2}})$ is connected. The same rule is applied in reverse to obtain $c^{(j_{2})}_{j_{1}}$. The descriptions and examples for dataset-specific rules are given in Appendix A.1.
@@
Motif Graph Construction.
Given a set of motifs, $V$, we describe our hierarchical abstraction over $V$. $G=(V,E)$ is a directed multigraph. Each $v\in V$ contains both the motif graph $g_{v}$ and $\{v_{r_{l}}\}$, denoting the possible “contexts” for attaching $g_{v}$ to another motif; that is, $\forall l,v_{r_{l}}\subseteq N(g_{v})$, with $N(g)$ denoting the set of atom nodes of graph $g$. $v_{R}:=\cup_{l}v_{r_{l}}$, and $\emptyset\neq v_{B}:=N(g_{v})\setminus v_{R}.$ Denoting $\sim$ to be the isomorphism relation, we construct $E$ by matching every pair of motifs $u,v$ and their contexts $(l_{1},l_{2})$ by finding corresponding subgraphs in $u_{B}$ and $v_{B}$ to match $u_{r_{l_{1}}}$ and $v_{r_{l_{2}}}$, as shown in Fig. 1 (c). Specifically, $(u,v,e_{l_{1},l_{2}})\in E\iff\exists b_{2}\subseteq u_{B},b_{1}\subseteq v_{B}$ such that:
@@
3.1 The Molecular Design Space as Derivations of a Context-Sensitive Grammar Over Motif Graph
We now define our context-sensitive grammar over $G$. We use the notations defined in the previous section to enumerate the set of production rules, $\mathcal{P}$, in our grammar. There is one initial rule $p_{v}\in\mathcal{P}$ for each motif $v$ in $G$, where the $LHS$ is $\mathcal{X}$, and the $RHS$ is the molecular graph $g_{v}$ with $u_{B}$ being the base atoms and $\{(u_{r_{l}})\}$ being the red atom sets that become “options” for attachment. Then, there is exactly one production rule $p_{u,v,l_{1},l_{2}}\in\mathcal{P}$ for each edge $(u,v,e_{l_{1},l_{2}})\in G$. This edge was attributed with $(u_{r_{l_{1}}},v_{r_{l_{2}}},b_{1},b_{2})$ during the construction of $G$. The application of the production rule then equates to attaching the fragment of $v$ to the fragment of $u$, at the attachment options keyed with $l_{1},l_{2}$. In the language of graph grammars, the context of this production rule is hence the molecular graph $g_{u}(u_{B}\cup u_{r_{l_{1}}})$, with the requirement that the matched atoms for $u_{r_{l_{1}}}$ are red. Applying this production rule replaces the matched atoms for $u_{r_{l_{1}}}$ within the $LHS$ by $g_{v}(N(g_{v})\setminus v_{r_{l_{2}}})$, where the red atom sets $\{v_{r_{l}}\mid v_{r_{l}}\cap v_{r_{l_{2}}}\neq\emptyset\}$ in $v$ are introduced as new options for attachment in the $RHS$. The random walk characterization arises out of the fact that if the $LHS$ molecule contains the context $g_{u}(u_{B}\cup u_{r_{l_{1}}})$, any edge $(u,v,e_{l_{1},l_{2}})\in E$ can be traversed, possibly including self-loops and parallel edges since $G$ is a directed multidigraph.
@@ 3.2 Molecules as Random Walks in the Design Space
Intuitively, our representation of a molecule $M$ captures a derivation in the above-defined context-sensitive grammar. While prior work has modeled such derivations in large and complex tree structures (e.g., with auxiliary nodes for partial derivations) (Guo et al., 2023a, b), we model it compactly in terms of a random walk over the bidirectionally connected subgraph $H_{M}=(V_{M},E_{M})$ of $G$ given by the fragmentation of $M$ [^3] Refer to Appendix B.3 for how and why we augment $G$ with duplicates of the same motif.; see Fig. 1 (a). Observe that $G$ is a strong prior for constraining the design space and sufficient for describing the molecular structure of $M$, but $H_{M}$ misses the global distribution of which it is a sample of.
Our learnable component models this distribution and, at the same time, captures the features that characterize a specific molecule in terms of a random walk. More specifically, our final molecule representation is a directed-acyclic multi-graph $\hat{H}_{M}=(V_{M},\hat{E}_{M},w_{M})$ that linearizes $H_{M}$ into a random walk such that (1) $\hat{E}_{M}\subseteq E_{M}$, (2) $\hat{H}_{M}$ remains connected, and (3) there is an Euler path [^4] In the case of monomers, the Euler path needs to be closed as monomers have the property of self-loops.. (i.e., each edge is used exactly once) $v_{0},v_{1},\ldots,v_{\ell}$ over $(V_{M},\hat{E}_{M})$ with $\hat{E}_{M}:=\cup_{i}\{(v_{i},v_{i+1})\}$; this path can be generated via a pre-order traversal that adds a reversed duplicate of the sub-trajectory when the stack contracts. The last component, $w_{M}$ is the sequence $w_{M}:=p_{0},p_{1},\dots,p_{\ell-1}$ of probabilities given by the random walk; that is, $p_{i}$ represents the probability with which the edge between $v_{i}$ and $v_{i+1}$ was traversed in the presence of all nodes visited thus far, as shown in Fig. 2. $w_{M}$ is parameterized by our learnable grammar, and $\hat{H}_{M}$ is explainable as a random walk of context-sensitive grammar rule applications.
Figure 2: Illustration of our generation procedure: (t=1) our learnable grammar parameterized by $\Phi$ samples a state transition $56\rightarrow 9$; (t=2) with the memory of having visited $\{56\}$, our grammar samples a state transition $\rightarrow 71$; (t=10) (bottom) our grammar samples a final transition $5$, which determines the molecular structure (top); our program’s notation is 56971[705]70:15:156$\rightarrow$ 9$\rightarrow$ 71[$\rightarrow$ 70$\rightarrow$ 5$\rightarrow$ 70:1% $\rightarrow$ 5:156 → 9 → 71 [ → 70 → 5 ] → 70 : 1 → 5 : 1
@@ 3.3.1 Parameter Estimation
where $D\in\mathbb{R}^{|V|\times|V|}$ is the in-degree matrix of $G$, $h(\cdot;\phi)$ is a memory-sensitive adjustment layer, and $c_{t}$ is a set-based memory of all nodes visited thus far. If $p^{(t)}$ is the current state of the random walk, the set-based memory, $c^{({t+1})}$, is updated as follows: $c^{(t+1)}\leftarrow\frac{t}{t+1}\cdot c^{(t)}+\frac{1}{t+1}\cdot p^{(t)}$. This set-based memory mechanism has precedents in graph theory literature. The learnable parameters are $\Phi=(W,\phi)$. Further motivation of the set-based memory mechanism is in Appendix C.3 and the full training algorithm can be found in Appendix D.
@@ 3.3.2 Training for a Downstream Task
Property Prediction Our grammar-induced molecular graph representation $\hat{H}_{i}$ allows for applying an off-the-shelf graph neural network $\mathcal{F}_{\Theta}$ to solve a given prediction task; in our evaluation, we used GIN (Xu et al., 2019). Given a property value $y^{(i)}\in\mathbb{R}$ with each molecule $M^{(i)}$, we apply a linear head $f_{\theta}$ and application-specific loss function $\mathcal{L}$ (e.g., MSE for regression or cross-entropy for classification).